# LD-MoLE: Learnable Dynamic Routing for Mixture of LoRA Experts

### v1

@@ 4.4 Sparsity control analysis
In this section, we evaluate how the proposed sparsity loss influences the expert pattern and impacts task performance on 5 datasets. To encourage sparsity, we set the target number of activated experts to $\leq 2$. Results show that the sparsity loss effectively reduces the overall number of activated experts. Increasing the sparsity alignment coefficient enforces a stronger constraint on the admissible range of $\lambda$ in Eq. 12. Moreover, Figure 4 illustrates the effect of applying the sparsity loss. Normally, more experts are activated in the lower layers, with activations gradually decreasing toward higher layers. Stronger regularization further suppresses higher-layer activations, while lower layers remain relatively dense. As shown in Table 3, the results emphasize a clear trade-off between sparsity and task performance: disabling the sparsity loss achieves the best average score, yet certain tasks benefit from reduced expert usage, suggesting that the optimal sparsity level is task-dependent. Our main contribution in this aspect is not simply confirming the sparsity and performance trade-off, but demonstrating that LD-MoLE makes sparsity both controllable and learnable within a dynamic routing framework. This enables the number of activated experts to be reduced in a principled way, while maintaining competitive performance.
@@ 4.5 Harder token need more experts
Dynamic routing allows the model to flexibly dedicate more capacity to tokens that require richer representations while conserving resources on more frequent or predictable ones. As shown in Figure 5, tokens that frequently appear during training (e.g., prompt- and context-related tokens) tend to activate fewer experts, effectively compressing their representations. In contrast, rarer or less familiar tokens activate a larger and more diverse set of experts, suggesting that tokens requiring greater modeling capacity benefit from richer expert combinations. This behavior is consistent with observations reported in ReMoE (Wang et al., 2025) for MoE pretraining. Overall, such adaptive routing enables the model to balance computation efficiently across tokens, allocating more resources to rare or informative ones.
@@ 5 More Ablation Studies
Additional exploration of LD-MoLE: In Appendix C.3, we present further experiments and discussions on our method, covering an alternative local MLP design as well as the effect of varying hidden dimensions in the shared MLP.
@@ 6 Conclusion
In this work, we introduce LD-MoLE, a learnable dynamic routing method for Mixture of LoRA Experts. Building on Sparsegen, our approach leverages a shared MLP to learn the sparsity parameter $\lambda$, enabling adaptive expert allocation across layers and tokens in a parameter-efficient manner. Comprehensive experiments show that LD-MoLE achieves the highest average scores on both the Llama-3.2-3B and Qwen3- 1.7B models compared to strong baselines, including TopK routing and ReLU-based routing, across a range of instruction-tuning and sequence classification tasks. For future research, we want to see how LD-MoLE performs in the pretraining stages of LLMs with its differentiability and controllable sparsity. Furthermore, integrating our dynamic routing framework with other PEFT methods or extending its applicability to new domains, such as multi-modal models, presents exciting opportunities for future exploration.

### v2

@@ 4.4 Sparsity control analysis
In this section, we evaluate how the proposed sparsity loss influences the expert pattern and impacts task performance on 5 datasets. To encourage sparsity, we set the target number of activated experts to $\leq 2$. Results show that the sparsity loss effectively reduces the overall number of activated experts. Increasing the sparsity alignment coefficient enforces a stronger constraint on the admissible range of $\lambda$ in Eq. 12. Moreover, Figure 4 illustrates the effect of applying the sparsity loss. Normally, more experts are activated in the lower layers, with activations gradually decreasing toward higher layers. Stronger regularization further suppresses higher-layer activations, while lower layers remain relatively dense. As shown in Table 3, the results emphasize a clear trade-off between sparsity and task performance: disabling the sparsity loss achieves the best average score, yet certain tasks benefit from reduced expert usage, suggesting that the optimal sparsity level is task-dependent. Our main contribution in this aspect is not simply confirming the sparsity and performance trade-off, but demonstrating that LD-MoLE makes sparsity both controllable and learnable within a dynamic routing framework. This enables the number of activated experts to be reduced in a principled way, while maintaining competitive performance. We also provide an additional experiment of computational analysis for this loss with respect to different hyperparameter $\beta$ in Appendix C.3.
@@ 4.5 Harder token need more experts
Dynamic routing allows the model to flexibly dedicate more capacity to tokens that require richer representations while conserving resources on more frequent or predictable ones. As shown in Figure 5, tokens that frequently appear during training (e.g., prompt- and context-related tokens) tend to activate fewer experts, effectively compressing their representations. In contrast, rarer or less familiar tokens activate a larger and more diverse set of experts, suggesting that tokens requiring greater modeling capacity benefit from richer expert combinations. This behavior is consistent with observations reported in ReMoE (ReLUMoE) for MoE pretraining. Overall, such adaptive routing enables the model to balance computation efficiently across tokens, allocating more resources to rare or informative ones.
@@ 5 More Ablation Studies
Additional exploration of LD-MoLE: In Appendix C.3, we present further experiments and discussions on our method. We provide an alternative local MLP design and analyze the impact of varying hidden dimensions in the shared MLP. In addition, we present experimental results on efficiency.
@@ 6 Conclusion
In this work, we introduce LD-MoLE, a learnable dynamic routing method for Mixture of LoRA Experts. Building on Sparsegen, our approach leverages a shared MLP to learn the sparsity parameter $\lambda$, enabling adaptive expert allocation across layers and tokens in a parameter-efficient manner. Comprehensive experiments show that LD-MoLE achieves the highest average scores on both the Llama-3.2-3B, Llama-3.2-8b and Qwen3- 1.7B models compared to strong baselines, including TopK routing and ReLU-based routing, across a range of instruction-tuning and sequence classification tasks. For future research, we want to see how LD-MoLE performs in the pretraining stages of LLMs with its differentiability and controllable sparsity. Furthermore, integrating our dynamic routing framework with other PEFT methods or extending its applicability to new domains, such as multi-modal models, presents exciting opportunities for future exploration.
7 Acknowledgements
The work of Yuan Zhuang and Fei Miao is partly supported by the National Science Foundation under Grants CNS-2047354 (CAREER). We also gratefully acknowledge the generous support of a Qualcomm gift fund and our collaborator Jonathan Petit.
@@ C.2 Expert Pattern During Training
Additional Experiment for the Sparsity Loss:
In this section, we compare the computational cost (in FLOPs) under different hyperparameter settings for our analytical loss function (Eq. 15). As shown in Table 6, the FLOP analysis further highlights the efficiency of the sparsity loss introduced in Section 3.2.2. As the sparsity-loss coefficient $\beta$ increases, the number of activated LoRA experts decreases, leading to a notable reduction in overall FLOPs. However, compared with conventional TopK and ReLU routing, the primary source of additional computation in our method arises from the shared MLP used to predict each token’s sparsity factor $\lambda$. To further mitigate this overhead, a promising direction is to augment an additional dimension into the gating projection for generating $\lambda$ and we leave it to the future exploration.
Qwen3-1.7B MoLA-8888 MoLA-2468 ReMoLE Ours($\beta=1.0$) Ours($\beta=0.1$) Ours($\beta=0.01$) Ours($\beta=0$)
MFLOPs 43 40 74 83 90 100 106
Table 6: Effective FLOPs (router + LoRA experts) across different $\beta$ parameter for the sparsity loss and different routing baseline. Backbone FLOPs are excluded since they are identical across methods.
@@
MMLU-Pro: It is designed to evaluate a model’s multi-domain knowledge understanding and complex reasoning abilities. Compared to MMLU, MMLU-Pro introduces more challenging, reasoning-focused questions and expands the answer choices from 4 to 10 options, while removing trivial and noisy items. Models using Chain-of-Thought reasoning perform better, reflecting the benchmark’s emphasis on deeper reasoning. The dataset contains 12,102 samples.